Mothballs are composed primarily of the hydrocarbon naphthalene (C10H8). When 1.025g of naphthalene burns in a bomb calorimeter, the temperature rises from 24.25∘C to 32.33∘C.
Find ΔrU for the combustion of naphthalene. The heat capacity of the calorimeter, determined in separate experiment, is 5.11kJ∘C−1.
1
Expert's answer
2016-02-10T00:01:19-0500
Combustion of naphthalene in the calorimetric bomb occurs at constant volume. Thus, change of system’s internal energy during the reaction will be equal to the amount of produced heat:
ΔrU = Q = CV·ΔT = 5.11·(32.33 – 24.25) = 41.3 kJ